OpenStudy (anonymous):

If f(x) = 6 - 5x, what is f(x) ^-1? -x-5/6 6+5/x 5x-6 6-x/5

OpenStudy (solomonzelman):

options are a little unclear.

OpenStudy (solomonzelman):

f(x)=6-5x (same as y=6-5x) -->x=(6-y)/5 -->f(y)=(6-y)/5 -->f^-1(x)=(6-x)/5

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

Anytime !